# CRS-Diff: Controllable Generative Remote Sensing Foundation Model

### v3

@@ Abstract
The emergence of generative models has revolutionized the field of remote sensing (RS) image generation. Despite generating high-quality images, existing methods are limited in relying mainly on text control conditions and thus don’t always generate images accurately and stablely. In this paper, we propose CRS-Diff, a new RS generative foundation framework specifically tailored for RS image generation, leveraging the inherent advantages of diffusion models while integrating more advanced control mechanisms. Specifically, CRS-Diff can simultaneously support text-condition, metadata-condition, and image-condition control inputs, thus enabling more precise control to refine the generation process. To effectively integrate multiple condition control information, we introduce a new conditional control mechanism to achieve multi-scale feature fusion, thus enhancing the guiding effect of control conditions. To our knowledge, CRS-Diff is the first multiple-condition controllable generative RS foundation model. Experimental results in single-condition and multiple-condition cases have demonstrated the superior ability of our CRS-Diff to generate RS images both quantitatively and qualitatively compared with previous methods. Additionally, our CRS-Diff can serve as a data engine that generates high-quality training data for downstream tasks, e.g., road extraction. The code is available at https://github.com/Sonettoo/CRS-Diff.
@@
Figure 1: (a) Comparison between natural image and remote sensing (RS) image. The image content is the Capital Museum of China, sourced from Google Maps and Google Street View, respectively. As can be seen, RS imagery differs significantly from traditional RGB imagery in resolution, coverage area, and information richness. (b) Comparison of the generation results between the two control modes. The upper image is the generation result guided solely by text, while the lower image is the result guided by both text and sketch. As can be seen, single text control condition fails to generate accurate image content while ”text + sketch” conditions can succeed.
@@ I Introduction
To address these issues, in this paper, we propose CRS-Diff, i.e. a controllable generative remote sensing foundation model. Specifically, a base diffusion model is first trained tailored for the RS domain based on the Stable Diffusion (SD) model [6], which is capable of converting high-precision textual descriptions into RS images as shown in Fig. 2 (a). Based on this, we integrated ControlNet [30] to include two additional control signals in the diffusion model for the controlled generation of RS images. These two control signals adjust global and local condition information of the image, including six additional image control conditions (semantic segmentation mask, roadmap, sketch, etc.) and textual conditions (prompt, content image, and metadata encoding) as shown in Fig. 2 (b). The optional combination of multiple conditions is controlled to ensure that the resulting RS images are visually realistic and accurately reflect specific geographic and temporal information. For the text condition, we concatenate directly with the original text encoding through an additional encoding step, leveraging the model’s natural control mechanism. For the image condition, we explore multiscale feature fusion to coordinate different control conditions and efficiently implement the bootstrapping of generative process noise maps, making our method flexible enough to combine any conditions for image generation, as shown in Fig. 2 (c).
@@
  • We propose a new controllable generative RS foundation model (CRS-Diff), which is a foundational framework specifically designed for RS image generation. Different from previous RS generative models, our CRS-Diff can simultaneously support more types of controllable conditions, i.e., text, metadata and image.
  • To effectively integrate multiple control information, we introduce a new conditional control mechanism to achieve multi-scale feature fusion to enhance the guiding effect of control conditions, thus broadening the image generation space. As far as we know, our CRS-Diff is the first multiple-condition controllable generative RS foundation model, which is capable of generating high-quality RS images that meet specific requirements under the guidance of composite conditions.
@@ III-A Text-to-Image generation
Textual information serves as the sole guiding factor in this process. Simultaneously, we employ the pre-trained CLIP ViT-L/14@336px model [6], fine-tuned on the RSICD RS image dataset, to amplify the effect of textual guidance. Under the original CFG framework, conditional noise prediction ${\epsilon}_{\theta}(z_{t},c)$ is solely dependent on the model’s processing of a given condition (e.g., text description). With the introduction of the CLIP bootstrap, this conditional prediction becomes further influenced by the similarity loss between the image and text as computed by the CLIP model. This implies that the calculation of ${\epsilon}_{\theta}(z_{t},c)$ incorporates considerations for better aligning the resulting image with the textual description.
@@ III-B Image Decoupling
Content: For content, the given image is considered as control information. Utilizing the Image Encoder in the pre-trained CLIP ViT-L/14@336px [41] model, the image is transformed into feature encoding, obtaining a global embedding. This approach offers more relevant guiding conditions than text descriptions alone.
@@ III-C Multi-conditional fusion
Text Conditional Fusion: We aim to establish a joint conditional bootstrapping mechanism that includes captions, content, and metadata, as illustrated in Fig. 4. The text description, as the main bootstrap condition, is represented as $y_{\text{t}}$ by obtaining a word embedding through a specialized CLIP text encoder. Considering that text input in T2I generation is a natural method, We map the metadata to an interval with a fixed range and encode it into a corresponding number of metadata encodings through a linear layer, with the result denoted as $y_{\text{m}}$. Additionally, we introduce an image encoder to encode the content image and then use a Feed-Forward Network (FFN) to connect the feature vectors with the prompt encoding symbols, represented as:
$y^{\prime}=\text{Concat}\left(y_{\text{t}},w_{\text{c}}\cdot\text{FFN}(y_{\text{c}}),w_{\text{m}}\cdot\text{MLP}(y_{\text{m}}\right)),$ (3)
@@ IV-B Implementation details
In the initial backbone model training phase, the following steps are taken: The RSICD dataset is fine-tuned over 10 epochs using the U-Net and the AdamW optimizer with a learning rate of $1\times 10^{-5}$. The input images are resized to $512\times 512$, with the model’s parameters totaling approximately 0.9 billion. For the sampling process, DDPM is utilized, setting the number of time steps to 100 and the classifier-free bootstrap scale to 7.5.
@@ IV-E Ablation Analysis
For ablation experiments, we explore improvements in the structure of the multi-conditional control network and the method of injecting control information. We perform ablation experiments on CRS-Diff and its variants, analyzing the impact of replacing the backbone model (ReB) and the feature fusion approach (FF) on the generation quality and control effectiveness, respectively. We constructed a baseline based on the underlying Multi-ControlNet and executed the alteration approach sequentially, reporting the evaluation metrics of the different models, as shown in Tab. V. The pre-training of the backbone model and the incorporation of the feature fusion approach significantly improve the generation effect, enabling the model to generate remote sensing images with higher information density and realize the fusion of various control information types, further enhancing the model’s control generation capability.
TABLE VI: Experimental results comparison of road detection in the setting of different training datasets. Both the real and synthetic datasets consist of 1000 images.
@@ IV-F Application for downstream road detection task
For the condition generation phase, we posit that the generated image should have a sufficiently high correlation with the control labels as conditions. We aim for the generated image to encapsulate as much information from the conditioned image as possible and to offer training data support for downstream tasks. Consequently, we incorporate experiments on the generated images to ascertain the pertinent performance of CRS-Diff. We consistently integrate synthetic dummy data into the training set of SGCN [50] for the road extraction task and assess it on the official test set. Tab. VI demonstrates the performance comparison of the SGCM method under different settings of training datasets. As can be seen, the synthetic training datasets can obtain almost the same performance as the real training dataset, which means that our CRS-Diff can simulate real images. By adding the synthetic dataset to the real dataset, the detection performance can be further significantly improved, indicating that the generated RS images conditioned by road are beneficial to the downstream road detection task. Additionally, Fig. 9 also visually compares the road extraction results under the three training conditions. The red boxes highlight areas where the differences are relatively large, revealing that the model trained with the synthetic dataset performs better in terms of continuity and completeness when dealing with a complex road network.
@@ V Conclusion
In this paper, we propose a new controllable generative RS foundation model (CRS-Diff). Developed from the diffusion model framework, CRS-Diff enables high-quality RS image generation. By integrating an optimized multi-conditional control mechanism, CRS-Diff can effectively synthesize multidimensional information, including text, metadata and images, guiding precise image generation and yielding highly accurate and controllable remote sensing images. This significantly broadens the control spectrum of the generation model, enhancing its adaptability to more complex application scenarios. Additionally, a comprehensive evaluation of existing multi-conditional generation models confirms CRS-Diff’s superior ability to generate remote sensing images under various conditions and its high controllability. This makes it suitable for a wide spectrum of use cases and enhances the performance of downstream tasks.

### v4

# CRS-Diff: Controllable Remote Sensing Image Generation with Diffusion Model

@@ Abstract
The emergence of generative models has revolutionized the field of remote sensing (RS) image generation. Despite generating high-quality images, existing methods are limited in relying mainly on text control conditions, and thus do not always generate images accurately and stably. In this paper, we propose CRS-Diff, a new RS generative framework specifically tailored for RS image generation, leveraging the inherent advantages of diffusion models while integrating more advanced control mechanisms. Specifically, CRS-Diff can simultaneously support text-condition, metadata-condition, and image-condition control inputs, thus enabling more precise control to refine the generation process. To effectively integrate multiple condition control information, we introduce a new conditional control mechanism to achieve multi-scale feature fusion, thus enhancing the guiding effect of control conditions. To our knowledge, CRS-Diff is the first multiple-condition controllable RS generative model. Experimental results in single-condition and multiple-condition cases have demonstrated the superior ability of our CRS-Diff to generate RS images both quantitatively and qualitatively compared with previous methods. Additionally, our CRS-Diff can serve as a data engine that generates high-quality training data for downstream tasks, e.g., road extraction. The code is available at https://github.com/Sonettoo/CRS-Diff.
@@
Figure 1: (a) Comparison between natural image and remote sensing (RS) image. The image content is the Capital Museum of China, sourced from Google Maps and Google Street View, respectively. As can be seen, RS imagery differs significantly from traditional RGB imagery in resolution, coverage area, and information richness. (b) Comparison of the generation results between the two control modes. The upper image is the generation result guided solely by text, while the lower image is the result guided by both text and sketch. As can be seen, the single text control condition fails to generate accurate image content while ”text + sketch” conditions can succeed.
@@ I Introduction
To address these issues, in this paper, we propose CRS-Diff, i.e. a controllable remote sensing generation model. Specifically, a base diffusion model is first trained tailored for the RS domain based on the Stable Diffusion (SD) model [6], which is capable of converting high-precision textual descriptions into RS images as shown in Fig. 2 (a). Based on this, we integrated ControlNet [30] to include two additional control signals in the diffusion model for the controlled generation of RS images. These two control signals adjust global and local condition information of the image, including six additional image control conditions (semantic segmentation mask, roadmap, sketch, etc.) and textual conditions (prompt, content image, and metadata encoding) as shown in Fig. 2 (b). The optional combination of multiple conditions is controlled to ensure that the resulting RS images are visually realistic and accurately reflect specific geographic and temporal information. For the text condition, we concatenate directly with the original text encoding through an additional encoding step, leveraging the model’s natural control mechanism. For the image condition, we explore multiscale feature fusion to coordinate different control conditions and efficiently implement the bootstrapping of generative process noise maps, making our method flexible enough to combine any conditions for image generation, as shown in Fig. 2 (c).
@@
  • We propose a new controllable RS generative model with diffusion models (CRS-Diff), which is a framework specifically designed for RS image generation. Different from previous RS generative models, our CRS-Diff can simultaneously support more types of controllable conditions, i.e., text, metadata and image.
  • To effectively integrate multiple control information, we introduce a new conditional control mechanism to achieve multi-scale feature fusion to enhance the guiding effect of control conditions, thus broadening the image generation space. As far as we know, our CRS-Diff is the first multiple-condition controllable RS generative model, which is capable of generating high-quality RS images that meet specific requirements under the guidance of composite conditions.
@@ III-A Text-to-Image generation
Textual information serves as the sole guiding factor in this process. Simultaneously, we employ the pre-trained CLIP ViT-L-14 model [6], fine-tuned on the RSICD RS image dataset, to amplify the effect of textual guidance. Under the original CFG framework, conditional noise prediction ${\epsilon}_{\theta}(z_{t},c)$ is solely dependent on the model’s processing of a given condition (e.g., text description). With the introduction of the CLIP bootstrap, this conditional prediction becomes further influenced by the similarity loss between the image and text as computed by the CLIP model. This implies that the calculation of ${\epsilon}_{\theta}(z_{t},c)$ incorporates considerations for better aligning the resulting image with the textual description.
@@ III-B Image Decoupling
Content: For content, the given image is considered as control information. Utilizing the Image Encoder in the pre-trained CLIP ViT-L-14 [41] model, the image is transformed into feature encoding, obtaining a global embedding. This approach offers more relevant guiding conditions than text descriptions alone.
@@ III-C Multi-conditional fusion
Text Conditional Fusion: We aim to establish a joint conditional bootstrapping mechanism that includes captions, content, and metadata, as illustrated in Fig. 4. The text description, as the main bootstrap condition, is represented as $y_{\text{t}}$ by obtaining a word embedding through a specialized CLIP text encoder. Different types of metadata are first mapped to values between 0 and 1 based on their value ranges, denoted as $\mathbf{m}=[m_{1},m_{2},\ldots,m_{n}]$, where $m_{i}$ denotes the $i$-th type of metadata. Subsequently, these normalized metadata values are encoded into vectors of uniform length using different Multi-Layer Perceptron (MLP) layers. These vectors are concatenated to form the metadata embedding $y_{\text{m}}$ as
$y_{\text{m}}=[\text{MLP}_{1}(m_{1});\text{MLP}_{2}(m_{2});\ldots;\text{MLP}_{n}(m_{n})],$ (3)
where $\text{MLP}_{i}$ denotes the MLP used for the $i$-th metadata.
Additionally, we introduce an image encoder to encode the content image and then use a Feed-Forward Network (FFN) to connect the feature vectors with the prompt encoding symbols, represented as:
$y^{\prime}=\text{Concat}\left(y_{\text{t}},w_{\text{c}}\cdot\text{FFN}(y_{\text{c}}),w_{\text{m}}\cdot y_{\text{m}}\right),$ (4)
@@ IV-B Implementation details
In the initial backbone model training phase, the following steps are taken: The RSICD dataset is fine-tuned over 10 epochs using the U-Net and the AdamW optimizer with a learning rate of $1\times 10^{-5}$. The input images are resized to $512\times 512$, with the model’s parameters totaling approximately 0.9 billion. For the sampling process, DDIM is utilized, setting the number of time steps to 100 and the classifier-free bootstrap scale to 7.5.
@@ IV-D2 Single-condition image generation
TABLE VI: Quantitative results comparisons among different CLIP Text encoder and training strategies.
Text encoder Parameters Fine-tuned IS $\uparrow$ FID $\downarrow$ CLIP Score $\uparrow$ Zero-Shot classification OA $\uparrow$
ViT-B-32 63M 15.73 83.21 19.93 55.46
16.22 73.94 19.96 63.26
ViT-L-14 123M 17.68 63.68 19.99 58.17
18.39 50.72 20.33 69.21
@@ IV-E Ablation Analysis
We explore improvements in the structure of the multi-conditional control network and the method of injecting control information. We perform ablation experiments on CRS-Diff and its variants, analyzing the impact of replacing the backbone model (ReB) and the feature fusion approach (FF) on the generation quality and control effectiveness, respectively. We constructed a baseline based on the underlying Multi-ControlNet and executed the alteration approach sequentially, reporting the evaluation metrics of the different models, as shown in Tab. V. The pre-training of the backbone model and the incorporation of the feature fusion approach significantly improve the generation effect, enabling the model to generate RS images with higher information density and realize the fusion of various control information types, further enhancing the model’s control generation capability.
TABLE VII: Experimental results under different control conditions for CLIP models ViT-L-14 and ViT-B-32. The best experimental results are in bold, and the second-best results are underlined.
CLIP Model Text Content image IS $\uparrow$ FID $\downarrow$ CLIP Score $\uparrow$
ViT-B-32 10.17 172.15 0.7379
10.64 95.98 0.7532
9.67 72.19 0.7671
ViT-L-14 10.99 135.76 0.7311
10.57 86.93 0.7573
9.62 72.30 0.7674
TABLE VIII: Quantitative results for different CLIP image encoders. The best experimental results are in bold, and the second-best results are underlined.
Image Encoder Parameters (M) IS $\uparrow$ FID $\downarrow$ CLIP Score $\uparrow$
ResNet-50 38 10.78 183.33 0.7239
ResNet-101 56 10.32 136.74 0.7383
ViT-B-32 87 10.15 119.30 0.7514
ViT-L-14 304 9.91 97.70 0.7528
Tab. VI presents a detailed ablation study to evaluate the impact of different versions of the CLIP model and the corresponding training strategy on the model generation capability in the single-text condition. We chose two versions of the CLIP text encoder for encoding the text condition in the text-to-image generation process, i.e., ViT-B-32 and ViT-L-14. We compared the effects of various parameter sizes and the impact of specific fine-tuning of the encoder on model performance. The results show that models with a greater number of parameters exhibit superior generation capabilities. Additionally, fine-tuning on RS images proves to be an effective method for enhancing performance.
We have conducted additional experiments to validate the positive impact of additional image control conditions on the quality of the generated images. We use a text encoder and an image encoder to process the text condition, the content image condition, and their combined conditions to generate the image, aiming to evaluate the quality of the generated images separately. The experimental results are shown in Tab. VII. As can be seen, this additional image condition information is beneficial to the quality of the generated images, reflecting on the FID and CLIP Score metrics. However, this image condition will reduce the diversity of the generated image, reflecting on the IS metric.
Tab. VIII describes the impact of different versions of the CLIP image encoder on model generation capability guided by the content image condition. We compared the results of image generation using four different CLIP image encoders and found that encoders based on the ViT architecture consistently achieved the best results in terms of FID and CLIP scores, which are standard metrics for evaluating image quality. The ViT-based encoders demonstrated a significant performance advantage over those based on the ResNet architecture. This suggests that the CLIP ViT-L-14 model currently in use possesses superior feature extraction capabilities.
TABLE IX: Experimental results comparison of road detection in the setting of different training datasets. Both the real and synthetic datasets consist of 1000 images.
@@ IV-F Application for downstream road detection task
For the condition generation phase, we posit that the generated image should have a sufficiently high correlation with the control labels as conditions. We aim for the generated image to encapsulate as much information from the conditioned image as possible and to offer training data support for downstream tasks. Consequently, we incorporate experiments on the generated images to ascertain the pertinent performance of CRS-Diff. We consistently integrate synthetic dummy data into the training set of SGCN [50] for the road extraction task and assess it on the official test set. Tab. IX demonstrates the performance comparison of the SGCM method under different settings of training datasets. As can be seen, synthetic training datasets can obtain almost the same performance as the real training dataset, which means that our CRS-Diff can simulate real images. By adding the synthetic dataset to the real dataset, the detection performance can be further significantly improved, indicating that the generated RS images conditioned on the road can promote the downstream road detection task. Besides, Fig. 9 also visually compares the road extraction results under the three training conditions. The red boxes highlight areas with relatively large differences, indicating that the model trained with the augmented dataset (Real + Synthetic) performs better in terms of continuity and completeness when dealing with complex road networks. Overall, the combination of real and synthetic data results in a more robust and generalized model, capable of handling more diverse and complex road structures. This blend of data sources not only increases the amount of training data but also introduces a wider range of scenarios, results in improved performance for road detection task.
@@ V Conclusion
In this paper, we propose a new controllable RS generative model with diffusion models (CRS-Diff). Developed from the diffusion model framework, CRS-Diff enables high-quality RS image generation. By integrating an optimized multi-conditional control mechanism, CRS-Diff can effectively synthesize multidimensional information, including text, metadata and images, guiding precise image generation and yielding highly accurate and controllable remote sensing images. This significantly broadens the control spectrum of the generation model, enhancing its adaptability to more complex application scenarios. Additionally, a comprehensive evaluation of existing multi-conditional generation models confirms CRS-Diff’s superior ability to generate remote sensing images under various conditions and its high controllability. This makes it suitable for a wide spectrum of use cases and enhances the performance of downstream tasks.